# A Hybrid Systems Model of Feedback Optimization for Linear Systems: Convergence and Robustness

### v4

@@ V-A States with Piecewise Constant Inputs
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\\ \cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$, (24)
where we define $t_{0}:=0$, the integer $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$ is the number of times the input to the LTI system has changed before hybrid time $(t,j)$, and $x(0,0)$ is a component of the initial condition $\phi(0,0)$.
@@
The model of $\mathcal{H}_{FO}$ applies piecewise constant inputs to the underlying LTI system. We will therefore integrate the underlying LTI dynamics over the intervals across which the input is constant and compute the difference between $x$ and $\tilde{x}$. For $p\in\mathbb{N}$, the input is equal to $u(t_{\bar{\alpha}(p)+p},\bar{\alpha}(p)+p)$ over intervals of the form $t_{\bar{\alpha}(p)+p},t_{\bar{\alpha}(p+1)+p+1}]$ and therefore we may write
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})+\int_{0}^{t}e^{A(t-\tau)}Bu(\tau)d\tau-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\\ =e^{At}x({0},{0})+\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\\ \cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$. (25)
@@ V-B Input Convergence
The following lemma relates successive iterates that are used to the compute the inputs to the LTI system. In it, we use
@@
Consider the hybrid system $\mathcal{H}_{FO}$ from (17) and consider an objective of the form of (8). Suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$. Let $\phi$ denote a maximal solution to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$ that satisfies (20). For any $(t,j)\in\textnormal{dom }\phi$, set $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$. Then, for any integer $p\in\{0,\ldots,P\}$ the state $z$ obeys
$\left\|z_{\alpha({p})}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha({p})+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha({p})+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ \leq q^{\frac{\alpha({p})-1}{2}}\left\|z_{1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|,$ (27)
@@
To ensure that $q\in\left(0,1\right)$, we use [27, Theorem 2.1.15], which shows that $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$ gives $q\in(0,1)$. Then iteratively applying (35) and taking the square root completes the proof. ∎
@@ V-C Complete Hybrid Convergence
$\mathcal{A}:={B_{r}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})}\times\mathcal{U}\times\mathcal{Y}\times\mathcal{U}\times[0,\tau_{c,\max}]\times[0,\tau_{g,comp}],$ (36)
where
$r=\frac{M\left\|B\right\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{\mathcal{U}}}}{\rho}(1+q^{\frac{\ell}{2}}),$ (37)
$M\geq 1$ is a constant,
@@
Consider the hybrid system $\mathcal{H}_{FO}$ from (17) and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (8), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$, where $Q_{u}$ is from (8) and $L:=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H)$. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$ that satisfies (20), for each $(t,j)\in\textnormal{dom }\phi$,
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +(M-1)\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big(1+q^{\frac{\ell}{2}}\big)\exp({-\rho}t),$ (39)
where $d_{\mathcal{U}}=\max_{u_{1},u_{2}\in\mathcal{U}}\|u_{1}-u_{2}\|$ is the diameter of the set $\mathcal{U}$, $\rho$ is from (38), $\ell\geq 1$ is from Assumption 2, $\mathcal{A}$ is from (36), $q:=1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\in\left(0,1\right)$, and $M\geq 1$ is a constant. In particular, each such solution satisfies
$\lim_{t+j\rightarrow\infty}\left\|\phi(t,j)\right\|_{\mathcal{A}}=0.$ (40)
@@
By definition of $\mathcal{A}$, only the state $x$ in $\phi$ affects the value of $\|\phi(t,j)\|_{\mathcal{A}}$. That is, we have
$\|\phi(t,j)\|_{\mathcal{A}}=\|x(t,j)\|_{B_{r}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})}$, (41)
and we therefore focus our analysis on $x$. We define $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$. Using (23), we observe that for each $p\in\{0,\ldots,P\}$ we have
@@
Next, using Lemma 2 and adding and subtracting $u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)$ we have
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau Bu^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$. (52)

### v5

@@ V-A States with Piecewise Constant Inputs
$x(t,j)-\tilde{x}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\\ \cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P})-\tilde{x}$, (22)
where we define $t_{0}:=0$, the integer $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$ is the number of times the input to the LTI system has changed before hybrid time $(t,j)$, and $x$ is the $x$ state component of the solution $\phi$.
@@
The model of $\mathcal{H}_{FO}$ applies piecewise constant inputs to the underlying LTI system. We will therefore integrate the underlying LTI dynamics over the intervals across which the input is constant and compute the difference between $x$ and $\tilde{x}$. For $p\in\mathbb{N}$, the input is equal to $u(t_{\bar{\alpha}(p)+p},\bar{\alpha}(p)+p)$ over intervals of the form $t_{\bar{\alpha}(p)+p},t_{\bar{\alpha}(p+1)+p+1}]$, over which our result follows.
@@ V-B Input Convergence
The following lemma relates successive iterates that are used to compute the inputs to the LTI system. In it, we use
@@
Consider the hybrid system $\mathcal{H}_{FO}$ from (15) and consider an objective of the form of (7). Suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$. Let $\phi$ denote a maximal solution to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$ that satisfies (18). For any $(t,j)\in\textnormal{dom }\phi$, set $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$. Then, for any integer $p\in\{0,\ldots,P\}$ the state component $z$ of $\phi$ obeys
$\left\|z_{\alpha({p})}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha({p})+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha({p})+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ \leq q^{\frac{\alpha({p})-1}{2}}\left\|z_{1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|,$ (24)
@@
To ensure that $q\in\left(0,1\right)$, we use [21, Theorem 2.1.15], which shows that $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$ gives $q\in(0,1)$. Then iteratively applying (32) and taking the square root completes the proof. ∎
@@ V-C Complete Hybrid Convergence
$\mathcal{A}:={B_{r}(\tilde{x}})\times\mathcal{U}\times\mathcal{Y}\times\mathcal{U}\times[0,\tau_{c,\max}]\times[0,\tau_{g,comp}],$ (33)
where $r=M\left\|B\right\|d_{\mathcal{U}}\rho^{-1}\big(2-\exp(-\rho\tau_{c,\min})+q^{\frac{\ell}{2}}\big)$, $M\geq 1$ is a constant, $q:=1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\in\left(0,1\right)$,
@@
Consider the hybrid system $\mathcal{H}_{FO}$ from (15) and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (7), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\big(0,\frac{2}{\lambda_{min}(Q_{u})+L}\big)$, where $Q_{u}$ is from (7) and $L:=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H)$. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$ that satisfies (18), for each $(t,j)\in\textnormal{dom }\phi$,
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +\frac{M^{2}\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(2-\exp(-\rho\tau_{c,\max})+q^{\frac{\ell}{2}}\Big)\exp({-\rho}t)\\ -\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1+q^{\frac{\ell}{2}}\exp({\rho}\tau_{c,\min})\Big)\exp({-\rho}t)$ (35)
where $d_{\mathcal{U}}=\max_{u_{1},u_{2}\in\mathcal{U}}\|u_{1}-u_{2}\|$ is the diameter of the set $\mathcal{U}$, $\rho$ is from (34), $\ell\geq 1$ is from Assumption 2, $\mathcal{A}$ is from (33), $q:=1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\in\left(0,1\right)$, and $M\geq 1$ is a constant. In particular, each such solution satisfies $\lim_{t+j\rightarrow\infty}\left\|\phi(t,j)\right\|_{\mathcal{A}}=0$.
@@
In this proof, we note that any empty sums evaluate to zero, and we take $\alpha(-1)=0$ for notational convenience. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$, the high-level steps of the proof are:
  1. 1.
    Bound the distance between the input $u$ and the optimal input $u^{*}$.
  2. 2.
    Express the difference $x(t,j{})-\tilde{x}$ in terms of integrals of the dynamics of $x(t,j)$.
  3. 3.
    Evaluate the resulting integral and simplify to bound $\|x(t,j{})-\tilde{x}\|$.
We begin with Step 1. By definition of $\mathcal{A}$, only the state component $x$ in $\phi$ affects the value of $\|\phi(t,j)\|_{\mathcal{A}}$. That is, we have
$\|\phi(t,j)\|_{\mathcal{A}}=\|x(t,j)\|_{B_{r}(\tilde{x})}$, (36)
and we therefore focus our analysis on $x$. We define $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$. Using (21), we observe that for each $p\in\{0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1},\ldots,P\}$ we have
@@
Onto Step 2, using Lemma 2 where $\tilde{x}$ is the optimal steady-state value and adding and subtracting $u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)$ we have
$x(t,j)-\tilde{x}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau Bu^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{x}$. (47)